OpenStudy (firejay5):

Show work!!! :D Medal may be rewarded! :O 5. (5 over y - 5) = (y over y - 5) - 1 6. (1 over 3x - 2) + (5 over x) = 0

OpenStudy (firejay5):

5. \[\frac{ 5 }{ y - 5 } = \frac{ y }{ y - 5 } - 1\] 6.\[\frac{ 1 }{ 3x - 2 } + \frac{ 5 }{ x } = 0\]

OpenStudy (anonymous):

@Firejay5 Are you ready for my explaination?? :)

OpenStudy (firejay5):

yea

OpenStudy (anonymous):

Ok so first I want to know if you know how to subtract fractions and numbers?

OpenStudy (firejay5):

1. I suck fraction so don't even try to ask me 2. I am partially good with numbers

OpenStudy (anonymous):

Ok that's ok I can help :) When we want to subtract fractions and numbers we need a common denominator. In this question lets work with the left side ok

OpenStudy (anonymous):

I mean the right side Sorry :P

OpenStudy (anonymous):

So lets begin :) This is what the equation is \[\frac{ y }{ y-5 }-\frac{ 1 }{ 1 }\] Do you agree??

OpenStudy (firejay5):

I guess

OpenStudy (anonymous):

Why aren't you confident??

OpenStudy (firejay5):

I thought it would be y - 5

OpenStudy (anonymous):

What do you mean? All I did to the right side of the equation was put 1 over 1 because all numbers are over 1. Are you still confused?

OpenStudy (firejay5):

see before you can add or subtract the denominator must be the same and all of the denominators are y - 5, so it should be 1 over y - 5

OpenStudy (anonymous):

You're on the right track but not quite. In order to have the same denominator you have to multiply both the numerator and denominator by \[y-5\] so it would be \[\frac{ y }{ y-5 }-\frac{ y-5 }{ y-5 }\] Does that make sense?

OpenStudy (firejay5):

see that's the point I was making

OpenStudy (anonymous):

Ok so can we proceed?? :P

OpenStudy (firejay5):

yes

OpenStudy (anonymous):

Ok so now we have the same denominator we can evaluate the numerators together \[\frac{ y-(y-5) }{ y-5 }\] does this make sense?

OpenStudy (firejay5):

yes

OpenStudy (anonymous):

Ok so now we see that there is a negative sign in front of the (y-5) so that negative will affect everything in the bracket. So... \[\frac{ y-y+5 }{ y-5 }\] Do you agree??

OpenStudy (firejay5):

yes

OpenStudy (anonymous):

Ok Can you take it from here then?

OpenStudy (firejay5):

so do we have so far?: \[\frac{ 5 }{ y - 5 } = \frac{ y - y + 5 }{ y - 5 }\]

OpenStudy (anonymous):

Yes but we are solving the right side in order to prove both sides. So finish the right side

OpenStudy (firejay5):

Do I cross multiply or what???

OpenStudy (anonymous):

No on the right side y-y+5 has two common terms which can be simplified

OpenStudy (firejay5):

what 2 common terms

OpenStudy (anonymous):

Y and Y so they cancel out

OpenStudy (firejay5):